# Convergence Analysis and Acceleration of Fictitious Play for General Mean-Field Games via the Best Response

### v1

@@ 2.1 A brief review of MFG and fictitious play
This paper aims to solve a finite-horizon Mean Field Game (MFG) system. The solution to a MFG is a pair of functions $(\rho,\phi)$, where $\rho(x,t)$ represents the population density and $\phi(x,t)$ denotes the value function of a representative player, both at position $x$ and time $t$. The value function $\phi$ solves a HJB equation involving $\rho$, indicating the influence of the population on the cost for a single player. Meanwhile, the population density $\rho$ solves a FP equation, where the drift is determined by $\phi$. To be precise, $(\rho,\phi)$ is a solution to a MFG if it solves
$-\partial_{t}\phi(x,t)-\nu\Delta\phi(x,t)+H(x,\nabla\phi(x,t))=f(x,\rho(\cdot,t)),$ $(x,t)\in Q_{T},$ (2)
$\partial_{t}\rho(x,t)-\nu\Delta\rho(x,t)-\nabla\cdot(\rho D_{p}H(x,\nabla\phi))(x,t)=0,$ $(x,t)\in Q_{T},$
$\phi(x,T)=f_{T}(x,\rho(\cdot,T)),\quad\rho(x,0)=\rho_{0}(x),$ $x\in\Omega.$
Here $\Omega\in\mathbb{R}^{d}$ is the space domain, $T>0$ is the terminal time and $Q_{T}:=\Omega\times[0,T]$ is the space-time joint domain. $\nu\geq 0$ is the physical viscosity. $H:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$ is the Hamiltonian. Let $\mathcal{P}(\Omega)$ be the set of probability distributions on $\Omega$. $\rho_{0}$ is a given initial distribution. $f:\Omega\times\mathcal{P}(\Omega)\to\mathbb{R}$ and $f_{T}:\Omega\times\mathcal{P}(\Omega)\to\mathbb{R}$ are the interaction cost and terminal cost. In addition, we impose homogeneous Neumann boundary conditions for $\rho$ and $\phi$, i.e.,
$\nabla\rho\cdot\mathbf{n}=0,\quad\nabla\phi\cdot\mathbf{n}=0,x\in\partial\Omega.$ (3)
For the existence and uniqueness of the solution to the MFG system, we refer to [18] and the references therein. In this paper, we restrict our focus on Hamiltonian $H(x,p)$ that are strictly convex in $p$ for any $x\in\Omega$ and denote the corresponding Lagrangian
$L(x,v):=\sup_{p}\left\{-\langle p,v\rangle-H(x,p)\right\}.$ (4)
We assume at least one solution exists to the MFG system (2) and we refer readers to [18, 27] and references therein for conditions for the existence to hold.
@@
Directly solving (2) is difficult because $\rho$ and $\phi$ are intertwined and the HJB is backward in time while the FP is forward in time. A natural idea to resolve this issue is to solve the HJB and FP equation alternatively. This leads to a fixed-point iteration. Precisely, during the $k$-th iteration, one calculates $\phi^{(k)}$ by solving the HJB equation with $\rho=\rho^{(k)}$, and then utilize $-D_{p}H(x,\nabla\phi^{(k)})$ to update $\rho^{(k+1)}$ through the FP equation. However, [6] shows that fixed-point iteration may jump between two states and fail to converge. Instead of using the latest population density as guidance, the fictitious play [19] uses a weighted average of history population densities. Specifically, when solving for the value function $\hat{\phi}^{(k+1)}$, an weighted average of history and current density, $\rho^{(k)}=\left(1-\delta_{k}\right)\rho^{(k-1)}+\delta_{k}\hat{\rho}^{(k)}~{}(0\leq\delta_{k}\leq 1)$, is utilized, rather than the current density $\hat{\rho}^{(k)}$. We summarize the procedure in Algorithm 1.
@@
$\partial_{t}\hat{\rho}^{(k)}-\nu\Delta\hat{\rho}^{(k)}-\nabla\cdot(\hat{\rho}^{(k)}D_{p}H(x,\nabla\hat{\phi}^{(k)}))=0,\quad\hat{\rho}^{(k)}(x,0)=\rho_{0}(x).$ (6)
@@
[19] shows $\delta_{k}=\frac{1}{k+1}$ leads to a convergent sequence. In section 2.2 of this paper, we give an explicit estimate of the convergence rate in terms of $\delta_{k}$ and show that other choices of weight $\delta_{k}$ can improve the convergence rate.
At the end of this part, we list two important concepts in MFG literature, monotonicity and potential games.
Definition 2.2 (Monotonicity).
$f:\Omega\times\mathcal{P}(\Omega)$ is monotone if for any $\rho,\rho^{\prime}\in\mathcal{P}(\Omega)$,
$\int_{\Omega}(f(x,\rho)-f(x,\rho^{\prime}))\mathrm{d}(\rho-\rho^{\prime})(x)\geq 0.$
$f$ is strictly monotone if the strict inequality holds for any $\rho\neq\rho^{\prime}$.
If both $f$ and $f_{T}$ are monotone, the MFG is called a monotone MFG. If the solution of a monotone MFG exists, then the solution is unique.
Definition 2.3 (Potential).
A MFG is a potential MFG if the interaction and terminal costs $f,f_{T}:\Omega\times\mathcal{P}(\Omega)\to\mathbb{R}$ derive from potentials, i.e. there exist $F,F_{T}:\mathcal{P}(\Omega)\to\mathbb{R}$ such that for any $\rho,\rho^{\prime}\in\mathcal{P}(\Omega)$,
$\lim_{s\to 0}\frac{F((1-s)\rho+s\rho^{\prime})-F(\rho)}{s}=\int_{\Omega}f(x,\rho)\mathrm{d}(\rho^{\prime}-\rho)(x),$
$\lim_{s\to 0}\frac{F_{T}((1-s)\rho+s\rho^{\prime})-F_{T}(\rho)}{s}=\int_{\Omega}f_{T}(x,\rho)\mathrm{d}(\rho^{\prime}-\rho)(x).$
When the game is potential and $\rho>0$, the MFG PDE system (2) is the optimality condition of the optimization problem
$\min_{\rho,m}$ $\int_{0}^{T}\int_{\Omega}\rho(x,t)L\left(x,\frac{m(x,t)}{\rho(x,t)}\right)\mathrm{d}x\mathrm{d}t+\int_{0}^{T}F(\rho(\cdot,t))\mathrm{d}t+F_{T}(\rho(\cdot,T))$ (8)

### v2

@@ 2.1 Second-order MFGs and monotonicity
Here $\rho(x,t)$ represents the population density and $\phi(x,t)$ denotes the value function of a representative player, both at position $x$ and time $t$. The value function $\phi$ solves a HJB equation involving $\rho$, indicating the influence of the population on the cost for a single player. Meanwhile, the population density $\rho$ solves a FP equation, where the drift is determined by $\phi$.
The two equations are coupled through the interaction cost $f$ and terminal cost $f_{T}$. We consider two common types of couplings $f$ and $f_{T}$ appear in the MFG literature. A nonlocal cost $f:\mathbb{T}^{d}\times\mathcal{P}(\mathbb{T}^{d})\to\mathbb{R}$ is used in the cases where players can influence each other even when they are not close. For example, $f(x,\rho):=\int_{\mathbb{T}^{d}}K(x,y)\mathrm{d}\rho(y)$, where $K:\mathbb{T}^{d}\times\mathbb{T}^{d}\to\mathbb{R}$ is a given kernel. Alternatively, one may consider local interaction costs, where $f:\mathbb{T}^{d}\times[0,+\infty)\to\mathbb{R}$ takes the form $f(x,\rho)=f(x,\rho(x))$ for absolutely continuous distributions $\rho$. A typical example is $f(x,\rho)=\log(\rho(x))$.
The notion of a weak solution in mean-field games (MFGs) varies depending on the setting, such as whether the cost is local or nonlocal, or whether the system is first-order or second-order. Correspondingly, the assumptions required for the existence of weak solutions also differ across these cases. For a comprehensive discussion, we refer the reader to Achdou2021MFGbook ; cardaliaguet2010notes ; Gomes2014survey and the references therein. For simplicity, we assume that the system under consideration admits at least one classical solution.
To ensure the uniqueness of the solution, monotonicity of interaction and terminal cost is necessary in general (Achdou2021MFGbook Remark 1.4). It also plays an important role in our convergence analysis.
Definition 2.1 (Monotonicity lasry2007mfg ).
A nonlocal cost $f:\mathbb{T}^{d}\times\mathcal{P}(\mathbb{T}^{d})$ is monotone if for any $\rho,\rho^{\prime}\in\mathcal{P}(\mathbb{T}^{d})$,
$\int_{\mathbb{T}^{d}}(f(x,\rho)-f(x,\rho^{\prime}))\mathrm{d}(\rho-\rho^{\prime})(x)\geq 0.$
$f$ is strictly monotone if the strict inequality holds for any $\rho\neq\rho^{\prime}$. For a local cost $f:\mathbb{T}^{d}\times[0,+\infty)$, we say $f$ is monotone if $f(x,\cdot)$ is nondecreasing and $f$ is strictly monotone if $f(x,\cdot)$ is increasing for any $x\in\mathbb{T}^{d}$.
If both $f$ and $f_{T}$ are monotone, we call the system (2) a monotone MFG.
If the solution to a monotone MFG exists and $H$ is strictly convex on $p$, then the solution is unique (Achdou2021MFGbook Theorem 1.4, Theorem 1.11 and Theorem 1.12).
2.2 Fictitious Play
@@
Directly solving (2) is difficult because $\rho$ and $\phi$ are intertwined and the HJB is backward in time while the FP is forward in time. A natural idea to resolve this issue is to solve the HJB and FP equations alternatively. This leads to a fixed-point iteration. Precisely, during the $k$-th iteration, one calculates $\phi^{(k+1)}$ by solving the HJB equation with $\rho=\rho^{(k)}$, and then utilize $-D_{p}H(x,\nabla\phi^{(k+1)})$ to update $\rho^{(k+1)}$ through the FP equation. However, Proposition 3.11 and the numerical experiment in Section 5.1 indicate that fixed-point iteration can be unstable, potentially oscillating between two states and failing to converge. Instead of using the latest population density as guidance, the fictitious play cardaliaguet2017ficplay uses a weighted average of history population densities. Specifically, when solving for the value function $\hat{\phi}^{(k+1)}$, an weighted average of history and current density, $\rho^{(k)}=\left(1-\delta_{k}\right)\rho^{(k-1)}+\delta_{k}\hat{\rho}^{(k)}~{}(0\leq\delta_{k}\leq 1)$, is utilized, rather than the current density $\hat{\rho}^{(k)}$. We summarize the procedure in Algorithm 1.
@@
[…]
@@
Remark 2.2 (Connection to the Folk Theorem wikipedia_folk_theorem ).
The use of a weighted average of historical population densities in fictitious play, rather than relying solely on the current density, emphasizes the role of history in shaping strategic decisions. This mirrors the insight of the Folk Theorem in repeated games, where long-term outcomes are sustained not only by current actions but also by the accumulated history of play. In both settings, past behavior influences future responses, enabling the emergence of equilibria that may not be supported in one-shot interactions. This connection highlights how memory and learning mechanisms can foster stable outcomes, even in decentralized or adaptive settings.
The following assumptions are enforced throughout the paper.
Assumption 2.3.
For the given initial distribution, Hamiltonian $H$, interaction cost $f$ and terminal cost $f_{T}$, we make the following assumptions:
  1. 1.
    The initial probability distribution $\rho_{0}$ is absolutely continuous with respect to the Lebesgue measure, and has a $C^{2}$ positive density, also denoted as $\rho_{0}$.
  2. 2.
    $H$ is $C^{2}$ in $x$, smooth and strictly convex in $p$, and satisfies
    $\lim_{|p|\to\infty}\frac{H(x,p)}{|p|}=+\infty,\quad\text{for any }x\in\mathbb{T}^{d}.$
  3. 3.
    $f$ satisfies one of the following assumptions:
    1. (a)
      $f$ is nonlocal, and $f(\cdot,\rho)\in C^{2}(\mathbb{T}^{d})$ for any $\rho\in\mathcal{P}(\mathbb{T}^{d})$.
    2. (b)
      $f$ is local, and $f\in C^{2}(\mathbb{T}^{d}\times[0,+\infty))$.
  4. 4.
    $f_{T}$ satisfies one of the following assumptions:
    1. (a)
      $f_{T}$ is nonlocal, and $f_{T}(\cdot,\rho)\in C^{3}(\mathbb{T}^{d})$ for any $\rho\in\mathcal{P}(\mathbb{T}^{d})$.
    2. (b)
      $f_{T}$ is local, and $f_{T}\in C^{3}(\mathbb{T}^{d}\times[0,+\infty))$.
Assumptions 2.3 do not guarantee the existence of a classical solution to the MFG system (2). However, standard results on semilinear parabolic equations ensure that they are sufficient to guarantee a unique classical solution at each step of the fictitious play algorithm, leading to the following lemma.
Lemma 2.4.
Under Assumption 2.3, if $\rho^{(0)}\in C^{2}(\overline{Q_{T}})$, $\rho^{(0)}(\cdot,t)\in\mathcal{P}(\mathbb{T}^{d})$, $\rho^{(0)}(\cdot,0)=\rho_{0}$, and $\rho^{(0)}>0$, then Algorithm 1 satisfies that for $k=1,2,\cdots$,
  1. 1.
    (3) has a unique classical solution $\hat{\phi}^{(k)}$.
  2. 2.
    (4) has a unique classical solution $\hat{\rho}^{(k)}$ and $\hat{\rho}^{(k)}>0$.
  3. 3.
    $\rho^{(k)}>0$.
2.3 Potential MFGs
In the literature, the assumption that the MFG has potential is key to proving the convergence of fictitious play.
Definition 2.5 (Potential).
A nonlocal cost $f:\mathbb{T}^{d}\times\mathcal{P}(\mathbb{T}^{d})\to\mathbb{R}$ or local cost $f:\mathbb{T}^{d}\times[0,+\infty)\to\mathbb{R}$ derive from potential if there exists $F:\mathcal{P}(\mathbb{T}^{d})\to\mathbb{R}$ such that for any $\rho,\rho^{\prime}\in\mathcal{P}(\mathbb{T}^{d})$,
$\lim_{s\to 0}\frac{F(\rho+s(\rho^{\prime}-\rho))-F(\rho)}{s}=\int_{\mathbb{T}^{d}}f(x,\rho)\mathrm{d}(\rho^{\prime}-\rho)(x).$
A MFG is a potential MFG if the interaction and terminal costs $f,f_{T}$ derive from potentials.
When the game is potential, the monotonicity of $f$ and $f_{T}$ is linked to the convexity of the potentials $F$ and $F_{T}$. Along with the convexity of $H$, lasry2007mfg points out that the MFG system (2) can be connected to the optimality condition of an optimization problem. The associated objective function then serves as a Lyapunov function in the convergence analysis.
To be precise, we first introduce a change of variable $(\rho,m)=(\rho,\rho v)$ and define the convex constraint set $C$
$C:=\left\{(\rho,m)\middle|\begin{aligned} &\rho\in C([0,T],L^{1}(\mathbb{T}^{d})\cap\mathcal{P}(\mathbb{T}^{d}));\\ &m\in L^{1}(\mathbb{T}^{d}\times[0,T];\mathbb{R}^{d});\\ &\rho(\cdot,0)=\rho_{0},\,\partial_{t}\rho-\nu\Delta\rho+\nabla\cdot m=0\end{aligned}\right\}.$ (6)
Here, the equations are satisfied in the sense of distributions. Let $L$ be the Lagrangian
$L(x,v):=\sup_{p}\left\{-\langle p,v\rangle-H(x,p)\right\},$ (7)
then $L$ is differentiable and strictly convex in $v$ by smoothness and strictly convexity of $H$ in $p$ (Assumption 2.3). We further define the dynamic cost $R$ as
$R({\rho},{m}):=\int_{0}^{T}\int_{\mathbb{T}^{d}}{\rho}(x,t)L\left(x,\frac{{m}(x,t)}{{\rho}(x,t)}\right)\mathrm{d}x\mathrm{d}t$ (8)
where when $\rho=0$, $L\left(x,\frac{m}{\rho}\right)$ is defined by convention
$L\left(x,\frac{m}{\rho}\right):=\begin{cases}0,&m=0,\\ +\infty,&m\neq 0.\end{cases}$ (9)
It is easy to verify that $R$ is jointly convex in $(\rho,m)\in C$. When $f,f_{T}$ derive from potentials $F,F_{T}$, denote the objective as
$J(\rho,m):=R(\rho,m)+\int_{0}^{T}F(\rho(\cdot,t))\mathrm{d}t+F_{T}(\rho(\cdot,T)).$ (10)
Then $J$ is jointly convex in $(\rho,m)$ if $F$ and $F_{T}$ are also convex. With the above notations, we have the following proposition for potential MFGs.
Proposition 2.6 (Potential MFGs).
Assume that $f$ and $f_{T}$ derive from potentials $F$ and $F_{T}$, and $F,F_{T}$ are convex under $L^{2}$ metric. If $(\rho^{*},\phi^{*})$ is of class $C^{2}(\mathbb{T}^{d}\times(0,T))$ with $\rho^{*}(x,0)=\rho_{0}(x)$, $\rho^{*}>0$, then $(\rho^{*},\phi^{*})$ is a solution to (2) if and only if $m^{*}(x,t):=-\rho^{*}D_{p}H(x,\nabla\phi^{*}(x,t))$ and
$J(\rho^{*},m^{*})=\min_{(\rho,m)\in C}J(\rho,m)$ (11)
Proof.
$(\rho^{*},\phi^{*})$ being a solution to (2) is sufficient due to the convexity of $F$ and $F_{T}$ and by the verification strategy used later in Lemma 3.1. We omit the details here.
For the necessity, $(\rho^{*},m^{*})$ is the minimizer and therefore $(\rho^{*},m^{*})\in C$. In addition, $m^{*}=-\rho^{*}D_{p}H(\cdot,\nabla\phi^{*})$. To show that $(\rho^{*},\phi^{*})$ solves the HJB equation in (2), we define
$\mathcal{L}(\rho,m;\phi^{*}):=J(\rho,m)-\int_{0}^{T}\int_{\mathbb{T}^{d}}\phi^{*}(x,t)(\partial_{t}\rho-\nu\Delta\rho+\nabla\cdot m)(x,t)\mathrm{d}x\mathrm{d}t.$ (12)
Then for any $(\rho,m)\in C$, $\mathcal{L}(\rho,m;\phi^{*})=J(\rho,m)$, and the optimality of $(\rho^{*},m^{*})$ gives
$\mathcal{L}(\rho^{*}+s(\rho-\rho^{*}),m^{*}+s(m-m^{*});\phi^{*})-\mathcal{L}(\rho^{*},m^{*};\phi^{*})\geq 0.$ (13)
Because $\rho^{*}>0$, we have that for any $(\rho,m)\in C$ and $s\in[0,1]$,
$R(\rho^{*}+s(\rho-\rho^{*}),m^{*}+s(m-m^{*}))-R(\rho^{*},m^{*})$ (14)
$=$ $s\int_{0}^{T}\int_{\mathbb{T}^{d}}(\rho-\rho^{*})\left(L\left(x,\frac{m^{*}}{\rho^{*}}\right)-\frac{m^{*}}{\rho^{*}}D_{v}L\left(x,\frac{m^{*}}{\rho^{*}}\right)\right)(x,t)\mathrm{d}x\mathrm{d}t$
$+s\int_{0}^{T}\int_{\mathbb{T}^{d}}(m-m^{*})\cdot D_{v}L\left(x,\frac{m^{*}}{\rho^{*}}\right)(x,t)\mathrm{d}x\mathrm{d}t+o(s)$
Therefore, we have that for any $(\rho,m)\in C$ and $s\in[0,1]$,
$\mathcal{L}(\rho^{*}+s(\rho-\rho^{*}),m^{*}+s(m-m^{*});\phi^{*})-\mathcal{L}(\rho^{*},m^{*};\phi^{*})$ (15)
$=$ $s\int_{0}^{T}\int_{\mathbb{T}^{d}}(\rho-\rho^{*})\left(\partial_{t}\phi^{*}+\nu\Delta\phi^{*}+L\left(x,\frac{m^{*}}{\rho^{*}}\right)-\frac{m^{*}}{\rho^{*}}D_{v}L\left(x,\frac{m^{*}}{\rho^{*}}\right)+f(x,\rho^{*}_{t})\right)(x,t)\mathrm{d}x\mathrm{d}t$
$+s\int_{0}^{T}\int_{\mathbb{T}^{d}}(m-m^{*})\left(D_{v}L\left(x,\frac{m^{*}}{\rho^{*}}\right)+\nabla\phi^{*}\right)(x,t)\mathrm{d}x\mathrm{d}t$
$+s\int_{\mathbb{T}^{d}}(\rho-\rho^{*})\Big{(}f_{T}(x,\rho^{*}_{T})-\phi^{*}\Big{)}(x,T)\mathrm{d}x+o(s)$
$=$ $-s\int_{0}^{T}\int_{\mathbb{T}^{d}}(\rho-\rho^{*})\Big{(}-\partial_{t}\phi^{*}-\nu\Delta\phi^{*}+H(x,\nabla\phi^{*})-f(x,\rho^{*}_{t})\Big{)}(x,t)\mathrm{d}x\mathrm{d}t$
$+s\int_{\mathbb{T}^{d}}(\rho-\rho^{*})\Big{(}f_{T}(x,\rho^{*}_{T})-\phi^{*}\Big{)}(x,T)\mathrm{d}x+o(s),$
where the first equality is by definition of $\mathcal{L}$ and potential, together with (14) and integration by part, and the second equality is because $\frac{m^{*}}{\rho^{*}}=-D_{p}H(\cdot,\nabla\phi^{*})$. Then by optimality (13), we conclude that $(\rho^{*},\phi^{*})$ solves the HJB equation.
@@ 2.3 Potential MFGs
For simplicity, we assume that $(\rho^{*},\phi^{*})\in C^{2}(\mathbb{T}^{d}\times(0,T))$ in Proposition 2.6. For more general conditions under which the MFG system (2) can be connected to a variational formulation, we refer the reader to Section 1.3.7.2 and Chapter 2 of Achdou2021MFGbook , as well as the references therein.
Since the potential formulation with monotone costs has a convex objective and linear constraint, many first-order optimization algorithms such as proximal gradient yu2024pgdeucmfg , primal-dual liu2021nonlocal and augmented Lagrangian benamou2015augmented ; benamou2017variational have been adapted to solve the problem besides fictitious play. While potential MFGs enjoy many desired properties, there are also limitations. For example, introducing $m=\rho v$ makes the Lipschitz constant of the objective function extremely large when $\rho$ is close to 0 and causes algorithms such as proximal gradient yu2024pgdeucmfg to be unstable. And there are MFGs not potential, for example $f(\cdot,\rho)=K*\rho$ where $K$ is not a symmetric kernel.
The convergence proofs in cardaliaguet2017ficplay ; Lavigne2023gcg-ficplay rely heavily on the decay of the functional $J(\rho,m)$, which is specific to potential games and does not extend to the general non-potential setting. One of the main contributions of this paper is to establish a convergence analysis of fictitious play for a broad class of MFGs without assuming the existence of a potential. In the potential case, cardaliaguet2017ficplay proves that setting $\delta_{k}=\frac{1}{k+1}$ yields convergence, while Lavigne2023gcg-ficplay connects fictitious play with the generalized conditional gradient method and analyzes how different choices of $\delta_{k}$ affect convergence rates. However, for non-potential games, the influence of $\delta_{k}$ on convergence has not been previously addressed. In Section 3, we provide an explicit convergence rate estimate in terms of $\delta_{k}$ and offer practical guidance on choosing $\delta_{k}$ to accelerate convergence in the non-potential setting.
3 Convergence and Acceleration of Fictitious Play
Given the fictitious play algorithm (Algorithm 1), several questions arise: Does it converge? Under what assumptions does convergence occur? And if it does, what is the rate of convergence? For potential MFGs, the objective function serves as a natural criterion for studying convergence. Based on this, cardaliaguet2017ficplay demonstrates the convergence of fictitious play for both first-order and second-order potential games, and Lavigne2023gcg-ficplay offers a sublinear convergence of classical fictitious play through the lens of the generalized conditional gradient. Lavigne2023gcg-ficplay also shows that a linear convergence rate can be achieved when the weight $\delta_{k}$ is chosen wisely.
In this section, we establish the linear convergence of fictitious play for non-potential MFGs. To that end, Section 3.1 introduces the key concept of the best response. This allows us to reformulate the fictitious play algorithm and define the best response gain as a measure of the distance to the Nash equilibrium. In Section 3.2, we present the assumptions and definitions underlying the convergence analysis, including directional strong convexity as a relaxation of global strong convexity, and discuss their validity. Section 3.3 then analyzes the stability of the best-response mapping, a central step in understanding the convergence behavior of fictitious play. Leveraging these stability properties, we show that the gain of the best response functions is a Lyapunov function and prove the convergence of fictitious play for non-potential MFGs; details are provided in Section 3.4. Section 3.5 presents the backtracking line search for optimizing the weight $\delta_{k}$ to accelerate the algorithm. Finally, Section 3.6 remarks on the convergence for potential MFGs, relating the gain of the best response to the dual gap in the generalized conditional gradient (Frank-Wolfe) algorithm.
3.1 The best response and the equilibrium formulation
When there is no ambiguity, we omit the input $x$ of $L,H,f$ and $f_{T}$.
We define the cost of $(\tilde{\rho},\tilde{m})$ given $\rho$ as
$J(\tilde{\rho},\tilde{m};\rho):=R(\tilde{\rho},\tilde{m})+\langle f(\rho),\tilde{\rho}\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}\rangle_{\mathbb{T}^{d}},$ (16)
where
$\langle f(\rho),\tilde{\rho}\rangle_{Q_{T}}:=\int_{0}^{T}\int_{\mathbb{T}^{d}}f(x,\rho_{t})\mathrm{d}\tilde{\rho}_{t}(x)\mathrm{d}t,\quad\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}\rangle_{\mathbb{T}^{d}}:=\int_{\mathbb{T}^{d}}f_{T}(x,\rho_{T})\mathrm{d}\tilde{\rho}_{T}(x),$
Here, the interaction cost and the terminal cost are linear in $\tilde{\rho}$. We first connect the fictitious play update (3)-(4) to an optimal control problem
Lemma 3.1.
For any $\rho\in\mathcal{P}(\mathbb{T}^{d})$, the optimal control problem
$\min_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho)$
admits a unique minimizer
$(\hat{\rho},\hat{m})=(\hat{\rho},-\hat{\rho}D_{p}H(\cdot,\nabla\hat{\phi})),$
where $(\hat{\rho},\hat{\phi})$ solves the decoupled system